OpenStudy (anonymous):

Which function is a linear function? 1 − 3x2 = -y y + 7 = 5x x3 + 4 = y 9(x2 − y) = 3 y − x3 = 8

OpenStudy (alexandervonhumboldt2):

a linear function is a function which has 1st power: common forms are y=mx+b Ax+By=c y-y_1=m*(x-x_1) linear functions are lines

OpenStudy (anonymous):

i still don't get it